OpenStudy (anonymous):

Find the missing number in the proportion. 3/8 = x/24

OpenStudy (anonymous):

\[\frac{3}{8}=\frac{x}{24}\iff x=\frac{3\times 24}{8}\]

OpenStudy (anonymous):

cancel first, multiply last

OpenStudy (anonymous):

Again, cross-multiply, you get 8x = 72. Divide both sides by 8, you get x = 9.